### v1

# DRAN: A Distribution and Relation Adaptive Network for Spatio-temporal Forecasting

Abstract
Accurate predictions of spatio-temporal systems’ states are crucial for tasks such as system management, control, and crisis prevention. However, the inherent time variance of spatio-temporal systems poses challenges to achieving accurate predictions whenever stationarity is not granted. To address non stationarity frameworks, we propose a Distribution and Relation Adaptive Network (DRAN) capable of dynamically adapting to relation and distribution changes over time. While temporal normalization and de-normalization are frequently used techniques to adapt to distribution shifts, this operation is not suitable for the spatio-temporal context as temporal normalization scales the time series of nodes and possibly disrupts the spatial relations among nodes. In order to address this problem, we develop a Spatial Factor Learner (SFL) module that enables the normalization and de-normalization process in spatio-temporal systems. To adapt to dynamic changes in spatial relationships among sensors, we propose a Dynamic-Static Fusion Learner (DSFL) module that effectively integrates features learned from both dynamic and static relations through an adaptive fusion ratio mechanism. Furthermore, we introduce a Stochastic Learner to capture the noisy components of spatio-temporal representations. Our approach outperforms state of the art methods in weather prediction and traffic flows forecasting tasks. Experimental results show that our SFL efficiently preserves spatial relationships across various temporal normalization operations. Visualizations of the learned dynamic and static relations demonstrate that DSFL can capture both local and distant relationships between nodes. Moreover, ablation studies confirm the effectiveness of each component.
Index Terms:
Spatio-temporal forecasting, graph neural network, distribution adaptation, adaptive network.I Introduction
Spatio-temporal systems, characterized by intricate spatial interactions among nodes and varying temporal dynamics, are prevalent in various fields such as physics [1], meteorology [2, 3, 4], power grids [5, 6] and transportation [7, 8, 9]. Real world spatio-temporal systems often entail a large number of nodes, with interactions between nodes varying over time [10, 11]. The complexity of these systems makes it challenging to make efficient decisions and manage future developments based on current conditions. Therefore, there is an urgent need for more accurate spatio-temporal prediction methods to support decision-making [12].
In spatio-temporal forecasting, historical time series data associated with nodes are used to predict future observations of the same nodes [13, 14]. Although numerous methods, particularly those based on deep learning have been developed [15, 16, 17, 18, 19], achieving accurate predictions remains challenging due to the time variance of the stochastic processes generating the data. This time variance can manifest in several ways, such as distribution shifts [20, 21], changes in spatial relationships among nodes [22], and changes in the nature of noise [23].
Normalization and de-normalization mechanisms are commonly employed to address distribution shifts in time series forecasting [24, 25, 26, 27]. Such techniques normalize input data to achieve consistent distributions and rebuild temporal distributions during de-normalization. Approaches such as Reversible Instance Normalization (RevIN) [24] and Dish-TS [27], which apply normalization and de-normalization on instance or temporal dimensions, have demonstrated their effectiveness in adapting to distribution shifts in time series forecasting. However, these methods may not be well-suited for spatio-temporal contexts. Methods that apply temporal normalization often scale the time series of nodes using different scaling factors. As a result, spatial message propagation occurs between nodes that have been scaled differently, which does not accurately reflect real-world message propagation processes. Moreover, instance-level normalization uses the mean and standard deviation of all nodes over the input time span, addressing only coarse spatial and temporal distribution shifts, which is inadequate for adapting to the complex distributions in spatio-temporal contexts. Therefore, it is necessary to develop approaches that achieve temporal normalization while preserving spatial patterns in spatio-temporal forecasting tasks.
To capture time-varying spatial relationships, some methods learn inherent static spatial relations from data [28, 29, 30], while others generate dynamic dependencies based on input windows [31, 32] . However, static relations remain fixed after training, limiting their adaptability, while dynamic relations derived from input data often exhibit instability in predictive performance. To address these limitations, some approaches jointly learn static and dynamic characteristics and fuse them with a weighted sum [33] or a fusion ratio constrained by loss functions [34], thus achieving more comprehensive spatial representation learning [35]. However, these approaches fail to account for the dynamic interplay between static and dynamic components, relying on fixed fusion ratios that result in insufficient spatio-temporal representations. Therefore, there is a pressing need for a network capable of adapting to both static and dynamic relations while fusing their features using adaptive rations.
Spatio-temporal systems inherently contain uncertainties due to data inaccuracies, noise, or sudden stochastic events within the systems. Given that these uncertainties impact prediction performance, various efforts have been made to quantify them. Some studies perform interval prediction to provide confidence regions for the forecasting horizon [36, 37]. Liang et al. [38] utilize latent variables to capture the uncertainties in air quality data, enhancing the accuracy of deterministic predictions.
In this paper, we propose a Distribution and Relation Adaptive Network (DRAN) that dynamically adapts to temporal variations and assesses uncertain components influenced by noise to achieve more comprehensive representation learning in spatio-temporal systems. More specifically, we develop a Spatial Factor Learner (SFL) that incorporates temporal normalization and de-normalization stages to mitigate the impact of distribution shifts while preserving spatial dependencies. To better accommodate spatio-temporal relations, we introduce a Dynamic-Static Fusion Learner (DSFL) module that fuses features learned from sample-specific dynamic and static relations using an adaptive fusion ratio mechanism. Additionally, we propose a Stochastic Learner based on a Variational Autoencoder (VAE) to estimate the noise. Experimental validations on weather and traffic systems to forecast temperature and traffic flow demonstrate that our model outperforms state-of-the-art methods. The SFL preserves the spatial distributions across various normalization methods. The learned dynamic and static relations capture non-overlapping local and distant node relations. Ablation studies further validate the effectiveness of each component. In summary, the novelty of our method is outlined as follows:
We introduce a SFL module that enables a normalization and de-normalization mechanism for distribution adaptation in a spatio-temporal context. The SFL module reverses temporal normalization while ensuring that spatial distributions closely match those of the forecasting horizons.
Different from previous spatio-temporal relations adaptive methods which fuse dynamic and static features with a fixed ratio [35, 39, 33, 34], we propose a DSFL module to adaptively combine features from both dynamic and static relations with a gating mechanism, enabling accommodating the changing relations of dynamic and static perspectives.
We propose a framework to learn the spatio-temporal representations from the deterministic and stochastic perspective. A VAE-based Stochastic Learner is introduced to learn the noisy components of representations.
The remainder of this paper is organized as follows. Section II introduces the current work on learning spatial and temporal relations and adapting distributions. Section III details our network architecture and the overall workflow. Section IV describes the experimental setups, including dataset descriptions, network configurations, training procedures, and baseline comparisons. Section V presents the numerical results for traffic flow and weather predictions, accompanied by an ablation study and visualizations of module impacts. To have a general view of the experimental results and the effectiveness of modules, we make a discussion in Section VI. Finally, we conclude the paper and outline future research directions in Section VII.
II Related Works
II-A Spatial and Temporal Relation Learning
Spatio-temporal forecasting methods are designed to capture both spatial and temporal relationships from static and dynamic perspectives. Static relations are typically learned from training datasets using trainable node embeddings and adjacency matrices to depict inter-node relationships [40, 17, 41]. Conversely, other studies [42, 43, 44] construct dynamic networks by generating sample-specific relations, which can be derived either directly from each input time series based on node similarity [45, 46] or through the use of meta-learners that generate dynamic parameters [47, 48]. Additionally, some methods refine the original relationships with input features [22, 49]. Furthermore, certain approaches simultaneously learn static and dynamic relations and integrate the gathered information. For instance, Fang et al. [34] decompose spatial relations into static and dynamic components, using a min-max learning paradigm with hyper-parameters to simultaneously learn both aspects. Similarly, Li et al. [33] employ two parallel Graph Convolutional Networks (GCNs) to learn static and dynamic information, which is then fused using a weighted sum. However, these methods typically fuse the static and dynamic features at a fixed ratio.
II-B Adaptation to Distribution Shifts
Normalization methods utilize statistic properties to normalize and de-normalize time series, adapting to distribution shifts in systems. Commonly, these methods utilize the mean and variance of input time series for data normalization. RevIN [24] employs a learnable affine transformation to align the means and standard deviations of inputs with those of outputs, facilitating distribution removal and reconstruction. Non-stationary Transformer [25] argues that existing stationary methods remove excessive information from time series, which hampers the model’s ability to learn temporal dependencies. To address this, it introduces De-stationary Attention modules that aim to balance this trade-off. Deep Adaptive Input Normalization (DAIN) [21] implements linear and gated layers to learn adaptive scaling and shifting factors for normalization. ST-Norm [50] proposes temporal and spatial normalization modules, which separately refines the high-frequency component and the local component of features. Additionally, U-Mixer [51] employs covariance analysis to correct stationarity for each feature. EAST-Net [52] generates sequence-specific network parameters to adapt dynamically to events.
III Methodology
III-A Problem Definition
The observed variables of node at time step can be referred to as a -dimensional vector . Each observation at time step is the result of a stochastic process, drawn from a conditioned distribution , where denotes the observations from nodes at time step in the spatio-temporal framework. The probability distribution varies over time and differs across nodes, indicating that distribution shifts occur when the observations across different time periods exhibit distinct distributions, i.e.,
where is a small threshold, is a distance function that estimates the discrepancy between distributions, and and refer to two different time steps. In this work, we utilize Gaussian kernel density estimation [53] for distribution estimation and use KL divergence-as distance function.
The goal of the spatio-temporal forecasting task is to develop a model that uses the historical observations of length of nodes to predict the future -step observations of nodes , where refers to time series of observed variables of all the nodes (). Due to the time variance in spatio-temporal data generation, learning a model that effectively handles shifting distributions is challenging. The overall workflow of DRAN and the structure of its modules are depicted in Fig. 1 and detailed in Algorithm 1.
III-B Distribution Adaptation
The distribution of historical time series often diverges from that of future time series due to time variance. Figure 2 (a) demonstrates the effectiveness of temporal normalization, which aligns the historical distribution more closely with the future distribution. This alignment facilitates easier learning of the projection from the input to the forecasting horizon. Consequently, temporal normalization is both an effective and necessary operation for spatio-temporal forecasting.
While previous works [25, 27] have addressed distribution shifts in multivariate time series forecasting, these methods are not suitable for time series with spatial relations. Direct application of these stationary strategies to spatio-temporal forecasting results in degraded performance. This is because they perform normalization within the time series of each node, leading to inconsistent scaling among nodes that are spatially interconnected. Therefore, we develop a framework that learns the temporal distribution shift on each node and preserves the spatial relations of nodes, thus maintaining the efficiency of spatial layers.
To learn the deterministic part of the inputs, We firstly filter the time series by removing high-frequency noise. We then normalize the input time series of each node using the temporal mean and standard deviation calculated for each input window.
where represents the normalized time series. To preserve essential temporal information, we incorporate temporal dependency learning through the de-stationary attention mechanism from the Non-stationary Transformer [25]. This mechanism utilizes de-stationary factors and , which are learned from a multilayer perceptron (MLP).
In this process, , , and represent the query, key, and value of attention, respectively, each derived from the projection of the input . The activation function is applied thereafter. To reverse the spatial patterns before inputting features into the spatial layers, we aim to de-normalize the features of nodes to preserve the spatial relations. Since the temporal representation is transformed by the De-stationary attention module, the input statistics and cannot be directly applied. As illustrated in Fig. 1 (c), we employ the SFL module to generate de-normalization factors and . In detail, we learn the node-wise features of the original input time series and the temporal representation using a 1-dimensional convolution on the temporal dimension and a Linear layer. Additionally, the statistics and are processed through a linear layer for feature dimension alignment.
Subsequently, we utilize a MLP to integrate node-wise features of the input, its temporal representation, and input data statistics to generate the de-normalization factors.
Finally, spatial factors and are applied to de-normalize .
where represents the de-normalized result of .
As illustrated in Fig. 2, our framework generates representations whose temporal and spatial distributions closely align with those of the forecasting horizon.
III-C Dynamic-Static Fusion Learner
Static neural networks with fixed parameters are inadequate for predicting dynamic systems. We consider that the relations of spatio-temporal systems consist of both static and dynamic components. Solely relying on input time series from different periods to establish spatio-temporal relations can lead to excessive fluctuations, thereby destabilizing prediction performance. Hence, it is essential to learn both dynamic and static information simultaneously and integrate them effectively.
To capture static information for a given task, we employ a trainable task-adaptive node embedding, akin to the approach used in the Adaptive Graph Convolutional Recurrent Network (AGCRN) [40]. Accounting for temporal variations within historical windows, we employ an adaptive node embedding to encode the static features of historical time series. Dynamic features are derived from the spatial similarity of each input data through spatial attention mechanisms.
where , , and represent the value, key, and query of spatial attention, respectively. represents the adjacency matrix capturing the learned dynamic spatial relations, and denotes the resulting dynamic features. Then, we extract static features through graph aggregation. In this process, we utilize a fully dense adjacency matrix, , derived from the adaptive node embedding , to depict the spatial relations.
where is the parameter matrix, and represents the features learned from static relations.
As the relationships between static and dynamic features change over time, we utilize a gating mechanism to integrate these features. The gate signal is generated to control the balance between dynamic and static features.
where conducts concatenation operation on the last dimension of dynamic features and static features , generating the concatenated features . We then generate the gate signal by applying a Sigmoid activation function to ; FC denotes the fully connected layer. Finally, we employ Equation (12) to fuse the dynamic and static features, producing the final deterministic features , which represent the final deterministic features learned from the input time series.
| Attributes | Weather | NYCBike1 | NYCBike2 | NYCTaxi | PeMS04 | PeMS08 |
| Start time | 1/1/2012 | 1/4/2014 | 1/7/2016 | 1/1/2015 | 1/1/2018 | 1/7/2016 |
| End time | 31/12/2022 | 30/9/2014 | 29/8/2016 | 1/3/2015 | 28/2/2018 | 31/8/2016 |
| Sample rate | 1 hour | 30 minutes | 30 minutes | 30 minutes | 5 minutes | 5 minutes |
| Training set | 5,623 | 3,023 | 1,912 | 1,912 | 10,181 | 10,700 |
| Testing set | 1,607 | 864 | 546 | 546 | 3,394 | 3,566 |
| Validation set | 803 | 431 | 274 | 274 | 3,394 | 3,566 |
| Node number | 263 | 128 | 200 | 200 | 307 | 170 |
| Feature number | 1 | 2 | 2 | 2 | 1 | 1 |
| Input length | 24 | 19 | 35 | 35 | 12 | 12 |
| Output length | 12 | 1 | 1 | 1 | 12 | 12 |
III-D Stochastic Learner
After obtaining the deterministic features, we utilize both backward and forward VAEs to generate the stochastic parts of the historical and forecasting time series.
Firstly, the deterministic features are fed into the latent layers to obtain the mean and standard deviation of , thereby capturing the distribution of the input data. Then, we sample the latent features from the distribution . These latent features are then processed through the reconstruction layer to map back to the stochastic components of the time series. The processes for the backward and forward Stochastic Learners are detailed below:
where and are the mean and standard deviation of the backward features, respectively, and and are those of the forward features. and represent the backward and forward latent layers, respectively, while and represent the backward and forward reconstruction layers. denotes the reconstruction of the historical time series, and refers to the stochastic parts of the forecasting time series. To ensure that the Stochastic Learner can effectively capture the input dynamic-related stochastic components, we impose constraints on the learned latent feature distribution and the reconstruction values.
The loss function consists of three components: prediction error , reconstruction error , and the distribution error computed using KL divergence .
where and are hyper-parameters that balance the importance of various loss functions, and represents the time series generated by the neural network. Both the prediction error and the reconstruction error are computed using the Mean Absolute Error (MAE).
| Task type | Methods | Task adaptive | Dynamic adaptive |
| Time series forecasting | Dual-stage Attention-based Recurrent Neural Network (DA-RNN) [54] | ✗ | ✗ |
| InfoTS [55] | ✓ | ✓ | |
| AutoTCL [56] | ✓ | ✓ | |
| Spatio-temoral forecasting | Temporal Graph Convolutional Network (TGCN)[28] | ✗ | ✗ |
| Spatio-Temporal Graph Convolutional Network (STGCN)[29] | ✗ | ✗ | |
| Graph Convolutional Gate Recurrent Unit (GCGRU) [30] | ✗ | ✗ | |
| Adaptive Graph Convolutional Recurrent Network (AGCRN)[40] | ✓ | ✗ | |
| Attention based Spatio-Temporal Graph Convolutional Networks (ASTGCN)[57] | ✗ | ✓ | |
| Diffusion Convolutional Recurrent Neural Network (DCRNN)[58] | ✗ | ✗ | |
| Spatio-Temporal Adaptive Embedding transformer (STAEformer)[32] | ✓ | ✓ | |
| Spatio-Temporal Self-Supervised Learning (ST-SSL)[22] | ✗ | ✓ | |
| Meta-Graph Convolutional Recurrent Network (MegaCRN) [59] | ✓ | ✓ | |
| Regularized Graph Structure Learning (RGSL) [17] | ✓ | ✗ | |
| Time-Enhanced Spatio-Temporal Attention Model (TESTAM) [60] | ✓ | ✓ | |
| Memory-based Drift Adaptation network (MemDA) [61] | ✓ | ✓ | |
| Ours | ✓ | ✓ |
After obtaining the deterministic component and the stochastic component of the time series, we input them into the decoder to map the features to the forecasting results. In this paper, the decoder comprises stacks of fully connected (FC) layers.
IV Experimental Setups
In this section, we detail our experimental setup, including the datasets used, the hyper-parameter settings for our networks, the baseline comparisons, and the training process.
IV-A Datasets
We conduct spatio-temporal forecasting tasks on weather and traffic systems to predict temperature and traffic flows.
Weather datasets.
For temperature forecasting, we use the ERA5 hourly dataset [62], originally with a resolution of , which we resample to . Our study focuses on an area between N to N latitude and E to E longitude, encompassing 263 nodes. The dataset covers the period from January 1, 2012, to December 31, 2022, and uses historical 24-hour temperature data to predict temperatures 12 hours ahead, with an input interval of 12 hours.
NYC datasets.
We use traffic flow datasets for bikes and taxis in New York, which have been preprocessed by Ji et al. [22]. These datasets are segmented into three categories: NYCBike1, NYCBike2, and NYCTaxi, recording inflows and outflows of city bikes and taxis every 30 minutes. The NYCBike1 dataset spans from April 1st, 2014, to September 30th, 2014. The NYCBike2 dataset covers the period from July 1st, 2016, to August 29th, 2016, while the NYCTaxi dataset ranges from January 1st, 2015, to March 1st, 2015. The input and output setups are consistent with those described in Ji et al. [22]. In NYCBike1, we predict the inflow and outflow of 128 grids 30 minutes ahead using historical records of 9.5 hours, with a sequence length of 19. In NYCBike2 and NYCTaxi datasets, we utilize historical time series of 17.5 hours, with a sequence length of 35. The number of grids in NYCBike2 and NYCTaxi datasets is 200.
PeMS04 and PeMS08 datasets.
The PeMS04 and PeMS08 datasets are subsets of the PeMS (PeMS Traffic Monitoring) dataset [63], which includes real-time traffic flow data collected from loop detectors on California highways. Specifically, PeMS04 contains traffic flow records from the San Francisco Bay Area, covering the period from January 1, 2018, to February 28, 2018. PeMS08 encompasses traffic data from July 1, 2016, to August 31, 2016. Both datasets are sampled at 5-minute intervals. In our forecasting task, we aim to predict traffic flow one hour ahead based on the past one-hour records. Both the input and output lengths for each prediction task are set to 12 time steps.
Application features of these datasets are shown in Table I.
| Weather | NYCBike1 | NYCBike2 | NYCTaxi | PeMS04 | PeMS08 | ||||||
| MAE | MAE | MAE | MAE | MAE | MAE | ||||||
| 0.30 | 0.765 | 1.50 | 5.115 | 1.00 | 4.903 | 1.50 | 10.961 | 0.50 | 19.283 | 0.50 | 13.884 |
| 0.50 | 0.719 | 3.50 | 5.073 | 1.50 | 4.875 | 2.50 | 11.039 | 0.70 | 18.614 | 0.70 | 13.879 |
| 0.80 | 0.705 | 5.50 | 5.158 | 2.50 | 4.914 | 3.50 | 10.864 | 0.90 | 18.585 | 0.90 | 13.933 |
| 1.00 | 0.660 | 7.50 | 5.057 | 3.50 | 4.827 | 4.50 | 11.074 | 1.00 | 18.561 | 1.00 | 13.726 |
| 3.00 | 0.766 | 9.50 | 5.159 | 4.50 | 4.847 | 5.50 | 11.030 | 2.00 | 18.615 | 1.20 | 13.696 |
| MAE | MAE | MAE | MAE | MAE | MAE | ||||||
| 0.10 | 0.670 | 0.10 | 5.390 | 0.10 | 5.130 | 0.10 | 11.264 | 0.10 | 32.290 | 0.10 | 20.790 |
| 0.50 | 0.665 | 0.50 | 5.269 | 0.50 | 4.910 | 0.50 | 10.758 | 0.50 | 25.774 | 0.50 | 13.818 |
| 1.00 | 0.664 | 1.00 | 5.173 | 1.00 | 4.866 | 1.00 | 11.136 | 1.00 | 18.585 | 1.00 | 13.717 |
| 5.00 | 0.658 | 5.00 | 4.982 | 5.00 | 5.025 | 5.00 | 11.114 | 5.00 | 18.545 | 5.00 | 13.822 |
| 10.00 | 0.672 | 10.00 | 5.303 | 10.00 | 4.938 | 10.00 | 11.161 | 10.00 | 18.387 | 10.00 | 13.696 |
| Selection () | (1.00, 5.00) | () | (7.50, 5.00) | () | (3.50, 1.00) | () | (3.50, 0.50) | () | (1.00,10.00) | () | (1.00, 1.00) |
IV-B Training Details
After the exploration stage, the averaged hyperparameters—assumed consistent across benchmarks—are presented as follows: The dimension of the adaptive node embedding is 80, consistent with STAEformer [32]. Two-layer MLPs with a hidden dimension of 64 are employed to generate and . In the SFL module, the Conv1d layers are configured with an input channel equal to the length of the lookback window , an output channel of 1, a kernel size of 3, and "circular padding" as defined in the PyTorch package. The Linear layers map the feature dimension to a hidden dimension of 64. For the DSFL module, both the de-stationary attention and spatial attention layers are set to 3. Each de-stationary attention module follows the parameter setup of STAEformer [32], with 4 attention heads and a feed-forward dimension of 256. In the DSFL module, the feature dimensions of and are both set to 160, and the number of attention heads for is also set to 4. In the Stochastic Learner, the latent layers include 3 Linear layers with ReLU activation functions, mapping the features to 64 dimensions. The reconstruction part of the Stochastic Learner consists of 3 Linear layers followed by ReLU activation functions, which remap the feature dimension from 64 to 160. The decoder comprises 2 Linear layers that fuse the deterministic and stochastic representations to produce the target feature outputs. The balance hyper-parameters and are fine-tuned experimentally to account for the stochastic nature and uncertainties of the datasets. Table III presents the variation in prediction error corresponding to different values of , with the value minimizing the error selected as optimal. Moreover, the training process is conducted on the Adam optimizer with a learning rate of 0.001 and a batch size of 32. The number of training epochs is set to 100. We split the datasets into training, validation and test set with a ratio shown in Table I. Numerical experiments for the methods are conducted using various random seeds from the set to obtain the average performance and standard deviation.
IV-C Baselines
We compare our method against several baseline approaches, including state-of-the-art multi-variable time series forecasting methods and spatio-temporal forecasting methods. Spatio-temporal forecasting techniques can be classified into two categories based on how they learn spatial relations: task-adaptive and dynamic-adaptive methods. Task-adaptive methods focus on learning static temporal and spatial relations from training datasets, which remain fixed during the testing phase. In contrast, dynamic-adaptive methods capture dynamically changing relations from input windows or by updating memory with incoming data. Details of the baseline methods are provided in Table II. In multi-variable time series forecasting, DA-RNN is a classic dual-stage attention-based recurrent neural network designed to capture long-term temporal dependencies. InfoTS and AutoTCL focus on enhancing time series representation learning through series augmentations and contrastive learning. InfoTS introduces a novel contrastive learning approach with information-aware augmentations that adaptively select optimal augmentations and a meta-learner network to learn from datasets. AutoTCL achieves unified and meaningful time series augmentations at both the dataset and instance levels, leveraging information theory to enhance representation quality. TGCN, STGCN, DCRNN and GCGRU utilize physical spatial relations as adjacency matrices and employ static neural networks for prediction. TGCN, DCRNN, and GCGRU use Graph Convolutional Networks (GCN) and Recurrent Neural Networks (RNN) to capture spatial and temporal features. STGCN combines temporal and graph convolutions to learn spatial and temporal dependencies. AGCRN utilizes learnable node embeddings to adapt spatial relations to tasks and node-adaptive parameters to capture specific attributes of each node. ASTGCN and STAEformer employ attention mechanisms to capture dynamic changes in input features. RGSL learns spatio-temporal dependencies from a predefined graph and learnable node embeddings. It dynamically fuses features from two graphs using an attention mechanism. Additionally, STAEformer employs learnable node embeddings and concatenates it with input time series to capture static features. MegaCRN utilizes node embeddings to learn the static relations and memory networks to dynamically match sample patterns with learned static features. ST-SSL does not learn static relations and only uses the adjacency matrix based on node distances as a prior graph. It fine-tunes the static graph using node similarities, effectively fusing dynamic and static features. TESTAM employs a mixture-of-experts model with three experts: one for temporal modeling, one for spatio-temporal modeling with a static graph, and one for spatio-temporal dependency modeling with a dynamic graph. We evaluate the performance of these models using two metrics: MAE and Mean Absolute Percentage Error (MAPE). The number of samples in the test datasets is denoted as . and denote the predicted and actual observations of spatio-temporal systems.
V Experimental Results
V-A Comparison Results
Numerical experiments, as shown in Tables IV and V, present the average performance and standard deviation of both the baseline methods and our DRAN model. Our DRAN model outperforms the baseline methods, demonstrating the best overall performance. Specifically, we observe a reduction in MAE of up to 7.8% in weather forecasting and 1.7% in traffic flow prediction compared to the baselines. DRAN achieves the best average performance on the Weather and NYC datasets. In terms of prediction stability, DRAN exhibits slightly larger standard deviations compared to other methods across experiments with different random seeds. This is likely due to the model’s attempt to learn noise components in the representation, which are randomly sampled from distributions. Our method performs less effectively on the MAPE metric but excels on the MAE metric. This suggests that the model prioritizes optimizing overall error rather than minimizing relative error in specific scenarios, such as scenarios with small target values. Consequently, while DRAN’s performance on the MAPE metric is slightly lower than that of other methods, its superior MAE results demonstrate effective overall error control. ST-SSL exhibits strong performance in one-step traffic flow prediction but is less effective on other datasets. This discrepancy may be due to ST-SSL’s emphasis on one-step ahead prediction, as its decoder is not optimized for multi-step forecasting. Additionally, MegaCRN performs well across both tasks, likely due to its task-adaptive node embeddings and the dynamic features learned from its meta memory networks. RGSL benefits from dynamically fusing an explicit prior graph with a learned implicit graph using attention mechanisms. STAEformer performs well on some datasets but poorly on others, which may be attributed to its non-adaptive fusion process. These results suggest that our proposed DRAN model is capable of adapting to various tasks and effectively capturing both static and dynamic features.
| Weather | NYCBike1 | NYCBike2 | NYCTaxi | |||||
| Temperature | Inflow | Outflow | Inflow | Outflow | Inflow | Outflow | ||
| DA-RNN [54] | MAE | 7.9511.036 | 14.1583.523 | 14.5653.383 | 13.4461.066 | 13.5131.382 | 72.88012.173 | 56.05610.753 |
| MAPE (%) | 2.7090.355 | 60.35912.515 | 62.95512.885 | 58.67813.627 | 58.80714.098 | 107.44121.603 | 98.41417.457 | |
| InfoTS [55] | MAE | 1.3430.031 | 6.4360.133 | 6.7970.138 | 6.6410.091 | 6.0870.096 | 14.9120.239 | 11.5450.225 |
| MAPE (%) | 0.4590.011 | 33.7240.553 | 34.9200.551 | 32.2310.634 | 30.1130.603 | 21.7700.442 | 20.9820.409 | |
| AutoTCL [56] | MAE | 1.1940.025 | 6.0020.037 | 6.4240.034 | 6.0110.060 | 5.5330.063 | 14.8430.162 | 11.3940.095 |
| MAPE (%) | 0.4080.008 | 28.0760.201 | 29.5730.164 | 29.5120.380 | 27.7660.409 | 22.0610.189 | 21.1400.136 | |
| STGCN [29] | MAE | 2.2781.103 | 17.1060.016 | 17.1770.216 | 32.98135.614 | 30.18028.265 | 78.62440.578 | 65.90733.232 |
| MAPE (%) | 0.7780.377 | 58.2081.926 | 58.7891.383 | 63.40219.134 | 62.23713.177 | 84.12430.571 | 75.09424.967 | |
| TGCN [28] | MAE | 1.7360.847 | 7.4120.052 | 7.6770.441 | 12.18510.559 | 10.7928.433 | 25.30711.321 | 21.7159.109 |
| MAPE (%) | 0.5920.287 | 34.6110.827 | 35.0851.704 | 37.5328.401 | 36.0457.883 | 44.53610.924 | 45.18110.584 | |
| MemDA [61] | MAE | 1.7860.071 | 6.7770.154 | 7.2460.100 | 6.7800.257 | 6.3440.286 | 20.9391.515 | 20.6311.268 |
| MAPE (%) | 0.6110.026 | 29.3440.543 | 30.3510.126 | 30.0990.342 | 28.5391.064 | 29.6513.243 | 34.9792.950 | |
| ASTGCN [57] | MAE | 1.5210.155 | 6.4810.352 | 6.6980.541 | 8.7235.180 | 7.5483.533 | 15.3895.711 | 12.5003.923 |
| MAPE (%) | 0.5200.054 | 30.7801.221 | 31.2641.976 | 28.6771.634 | 27.1992.117 | 24.7310.916 | 24.4431.578 | |
| TESTAM [60] | MAE | 1.4811.203 | 6.3310.736 | 6.8260.891 | 6.5580.941 | 6.3081.003 | 22.6863.816 | 21.4504.195 |
| MAPE (%) | 0.5070.415 | 30.4182.571 | 31.6282.949 | 30.1603.249 | 29.7054.047 | 32.1435.566 | 37.2377.208 | |
| AGCRN [40] | MAE | 4.8522.311 | 6.3220.847 | 6.5250.510 | 10.8645.159 | 9.8893.836 | 19.7868.580 | 16.4326.402 |
| MAPE (%) | 1.6690.798 | 30.0491.915 | 30.5751.184 | 34.9993.415 | 34.0323.307 | 33.3076.145 | 32.1815.601 | |
| GCGRU [30] | MAE | 1.0120.041 | 5.4570.063 | 5.6310.290 | 7.0363.489 | 6.2502.526 | 12.1532.649 | 10.2771.340 |
| MAPE (%) | 0.3460.015 | 26.8670.700 | 27.2921.476 | 24.3372.990 | 23.7112.231 | 22.3867.043 | 22.8007.607 | |
| DCRNN [58] | MAE | 0.9840.028 | 5.3740.039 | 5.5570.305 | 6.9403.500 | 6.1532.513 | 12.0973.639 | 9.8332.314 |
| MAPE (%) | 0.3360.010 | 26.7720.850 | 27.1651.507 | 24.0703.034 | 23.5652.251 | 21.1823.214 | 21.5983.360 | |
| STAEformer [32] | MAE | 3.7282.367 | 5.1680.029 | 5.4750.028 | 5.4530.100 | 5.1120.147 | 12.2620.237 | 9.8240.109 |
| MAPE (%) | 1.2840.819 | 25.8290.331 | 26.8890.242 | 25.7740.429 | 24.8400.447 | 17.8890.777 | 18.2030.684 | |
| MegaCRN [59] | MAE | 0.9520.027 | 5.0420.016 | 5.3570.043 | 6.6023.125 | 5.8782.164 | 12.2060.083 | 9.7400.074 |
| MAPE (%) | 0.3250.010 | 25.3290.167 | 26.2750.213 | 23.4233.445 | 22.7102.775 | 18.0310.582 | 17.9720.294 | |
| RGSL [17] | MAE | 0.7270.003 | 5.1490.140 | 5.3350.200 | 6.9213.513 | 6.0822.518 | 13.9451.775 | 11.8592.918 |
| MAPE (%) | 0.2480.001 | 25.6120.183 | 26.1100.924 | 24.1862.957 | 23.2882.249 | 27.14817.848 | 27.37117.905 | |
| ST-SSL [22] | MAE | 1.3940.040 | 5.1350.024 | 5.2650.023 | 5.0420.029 | 4.7140.023 | 12.0100.481 | 9.7900.101 |
| MAPE (%) | 0.4750.013 | 25.4300.295 | 24.6050.265 | 22.6330.112 | 21.8130.808 | 16.3830.100 | 16.8550.228 | |
| Ours | MAE | 0.6720.007 | 4.8820.032 | 5.1760.054 | 5.0080.056 | 4.6530.036 | 11.929 0.054 | 9.5390.054 |
| MAPE (%) | 0.2290.002 | 23.5530.169 | 24.6070.497 | 22.3840.211 | 21.4290.208 | 16.3380.409 | 16.6660.201 | |
Results with bold are the overall best performance, and shading results have the suboptimal performance.
Furthermore, we display the prediction results of our DRAN and the sub-optimal methods. As is shown in Fig. 3, by comparing the prediction errors of our DRAN, RGSL, and MegaCRN, we find that while the numerical metrics of these methods are very close, the distribution of prediction errors is different. The prediction results of our method demonstrate a more stable performance across all spatial regions, whereas other methods exhibit significantly larger errors in certain regions. This indicates that our method is more stable in prediction and better adapts to nodes with complex dynamic changes. As shown in Fig. 4, we provide two cases as examples to visualize the prediction results. We can see that both methods generate predictions that are similar to the ground truth. However, when comparing spatial prediction errors, DRAN shows fewer grids with large prediction errors. Additionally, we present the predicted time series of nodes in Fig. 5. In the weather dataset, both RGSL and our method capture the overall temporal trends of nodes and perform well in the more regular periodic variations, though they lack accuracy in some extreme values. This may be due to an insufficient ability to capture abrupt changes in the time series. In Fig. 5 (c) and (d), DRAN demonstrates superior performance in predicting the sudden decrease in traffic flow.
To balance computational cost and prediction accuracy, we compare inference times in Table VI, where methods are listed in descending order of inference time. The comparison is conducted using input data of the same size, repeated 100 times to compute the average inference time for a batch of NYCTaxi data. All experiments are performed on an Nvidia RTX3090 GPU. While DA-RNN achieves the fastest inference speed, it lacks sufficient prediction accuracy. DRAN delivers the best prediction accuracy with a moderate latency at inference time, comparable to ST-SSL and STAEformer.
V-B The Preservation of Spatial Distribution
To evaluate whether the SFL module preserves spatial distributions across various tasks or not, we analyze the spatial distributions of representations before SFL (), after SFL (), and at the forecasting horizon (). The objective is to determine whether the distributions of representations after SFL are closer to those at the forecasting horizon. For each forecasting task, we randomly generate indices and select samples from , , and . We then use Gaussian Kernel Density Estimation to assess the distributions of these representations. As visualized in Fig. 6, the results demonstrate the SFL module’s ability to preserve the spatial distribution of nodes. The primary goal of SFL is to align the spatial distribution of learned representations with that of the forecasting horizon, thereby enhancing accuracy in capturing spatial and temporal distribution changes. In Fig. 6, significant discrepancies between the spatial distribution of representations before SFL and the final horizon are observed. The SFL module effectively reduces these variations in the spatial distribution of learned representations, thereby facilitating the learning process to map spatially preserved representations to predictions.
Furthermore, we replace the temporal normalization modules to verify the effectiveness of SFL across various temporal normalization operations. We assessed SFL’s ability to preserve spatial distribution when combined with different temporal normalization modules: DAIN, Dish-TS, the Non-stationary Transformer and ST-norm. DAIN utilizes MLPs and a gate mechanism to learn the adaptive mean and standard deviation of input time series. The Non-stationary Transformer normalizes the time series and employs scaling factors to prevent removing excessive temporal information within the attention module. Dish-TS normalizes and de-normalizes the lookback horizon windows using different learned means and standard deviations. We applied temporal normalization after the frequency cutting operation, with SFL using the mean and standard deviation of the lookback windows to generate spatial factors. Given that RevIN performs normalization on feature dimensions, we conducted experiments applying only feature normalization to investigate whether this coarse-grained approach, which simultaneously normalizes both spatial and temporal distributions, is sufficient for spatio-temporal forecasting tasks. ST-norm applies spatial and temporal normalization to the inputs, enabling the model to better capture high-frequency spatial features and local temporal features. We compare three scenarios: applying T-norm only, applying ST-norm, and combining T-norm with the SFL module.
As shown in Table VII, SFL improves the performance of various temporal normalization methods, demonstrating its effectiveness as a general module for spatial distribution preservation. The models incorporating temporal operations and SFL outperform the model with RevIN. This finding suggests that normalization on feature dimensions alone is too coarse-grained and may not be adequate for distribution adaptation in spatio-temporal tasks. In ST-norm, the combination of spatial and temporal normalization improves performance. However, combining T-norm with the SFL module results in greater accuracy improvement, suggesting that the rescaling spatial distributions contribute more significantly to imitating propagation dynamics than normalization alone. Therefore, SFL after temporal normalization is an effective approach for distribution adaptation compared with instance-level and spatial and temporal level normalization.
V-C Dynamic and Static Relations Learning
In Fig. 7, we explore the adaptive process of our model by depicting the dynamic and static adjacency matrices within the DSFL module. Specifically, we showcase these matrices for the first input time step as an example. The dynamic adjacency matrix is derived from the similarity of time series between nodes according to Equation 6, while the static adjacency matrix is obtained by learning the static relations according to Equation 8. Fig. 7 illustrates the spatial relations between nodes of Weather dataset, with darker colors indicating stronger relationships. The dynamic adjacency matrix highlights the strength of relationships between nodes with similar signal patterns, whereas the static adjacency matrix focuses on the signals of individual nodes and some distributed nodes within the network. The differing concentrations of dynamic and static perspectives allow the model to learn features from various aspects.
To clarify the differences between the learned dynamic and static relations for specific nodes, we select node from various locations and visualize the relationships between the selected node and other nodes, represented as and . As shown in Fig. 8, subfigures (a), (b), and (c) depict the dynamic relations between target nodes and other nodes in the Weather dataset, while subfigures (d), (e), and (f) illustrate the static relations. The dynamic relations learned by DSFL concentrate around the target nodes, highlighting the significance of local connections. In contrast, the static relations reflect interactions between target nodes and distant nodes, indicating that the static adjacency matrix captures non-local relationships. This demonstrates that DSFL effectively learns comprehensive and non-overlapping spatial relations.
| PeMS04 | PeMS08 | ||
| Flow | Flow | ||
| DA-RNN [54] | MAE | 130.38427.184 | 110.05619.332 |
| MAPE (%) | 178.98325.358 | 100.13928.297 | |
| InfoTS [55] | MAE | 25.8510.510 | 24.0061.225 |
| MAPE (%) | 19.5560.464 | 14.6820.666 | |
| AutoTCL [56] | MAE | 23.8140.048 | 17.3550.200 |
| MAPE (%) | 20.8790.098 | 13.0060.121 | |
| TGCN [28] | MAE | 34.78590.2043 | 33.6049.765 |
| MAPE (%) | 27.9720.754 | 26.1729.541 | |
| GCGRU [30] | MAE | 25.83360.0399 | 22.2658.494 |
| MAPE (%) | 17.7880.442 | 14.3018.571 | |
| STGCN [29] | MAE | 25.30172.3161 | 25.33810.825 |
| MAPE (%) | 24.3218.603 | 15.4047.083 | |
| DCRNN [58] | MAE | 24.91172.0638 | 20.8530.062 |
| MAPE (%) | 17.7200.973 | 11.8640.094 | |
| ASTGCN [57] | MAE | 23.56480.9421 | 20.3080.967 |
| MAPE (%) | 16.8131.004 | 11.3030.552 | |
| ST-SSL [22] | MAE | 23.1461.074 | 18.9890.713 |
| MAPE (%) | 14.4130.697 | 10.7980.332 | |
| MemDA [61] | MAE | 20.0370.150 | 16.3700.207 |
| MAPE (%) | 11.9690.100 | 9.3420.243 | |
| RGSL [17] | MAE | 19.544+-0.2571 | 17.4523.102 |
| MAPE (%) | 13.8620.242 | 9.1860.182 | |
| TESTAM [60] | MAE | 19.3310.481 | 15.7570.360 |
| MAPE (%) | 12.0980.419 | 9.0830.215 | |
| AGCRN [40] | MAE | 19.32910.3053 | 17.7902.210 |
| MAPE (%) | 12.9370.035 | 10.3241.078 | |
| MegaCRN [59] | MAE | 18.8580.0413 | 15.5970.245 |
| MAPE (%) | 12.8080.098 | 8.8340.096 | |
| STAEformer [32] | MAE | 18.2410.082 | 13.5380.039 |
| MAPE (%) | 12.0640.071 | 8.8580.017 | |
| Ours | MAE | 18.3750.087 | 13.6900.085 |
| MAPE (%) | 12.0260.419 | 8.9870.057 | |
Results with bold are the overall best performance, and shading results have the suboptimal performance.
| Methods | Time (s) | MAE |
| GConvGRU [30] | 1.383 | 12.022 |
| DCRNN [58] | 0.967 | 11.924 |
| TGCN [28] | 0.469 | 28.475 |
| STGCN [29] | 0.343 | 12.286 |
| RGSL [17] | 0.227 | 11.916 |
| InfoTS [55] | 0.179 | 13.228 |
| MegaRCN [59] | 0.166 | 10.970 |
| AGCRN [40] | 0.134 | 18.338 |
| ASTGCN [57] | 0.114 | 14.422 |
| TESTAM [60] | 0.075 | 24.303 |
| DRAN (Ours) | 0.075 | 10.737 |
| ST-SSL [22] | 0.073 | 10.996 |
| STAEformer [32] | 0.065 | 10.810 |
| AutoTCL [56] | 0.052 | 13.119 |
| MemDA [61] | 0.047 | 20.521 |
| DA-RNN [54] | 0.038 | 64.468 |
Experiments are conducted on a batch of NYCTaxi date which contains 8 samples. Results with bold are the fastest method and the method with best prediction performance.
| Weather | NYCBike1 | NYCBike2 | NYCTaxi | |||||
| Temperature | Inflow | Outflow | Inflow | Outflow | Inflow | Outflow | ||
| +RevIN | MAE | 0.7320.004 | 5.0310.064 | 5.3410.094 | 5.1920.068 | 4.8310.060 | 12.3010.217 | 9.7140.168 |
| +DAIN | MAE | 1.0350.006 | 5.2120.088 | 5.5080.084 | 5.6700.420 | 5.2580.290 | 13.7470.174 | 10.8480.088 |
| +DAIN+SFL | MAE | 0.6630.003 | 4.9420.050 | 5.2290.028 | 5.2680.135 | 4.8990.159 | 12.3030.147 | 9.8280.171 |
| +Non-st | MAE | 0.7380.006 | 5.0960.109 | 5.3680.089 | 5.2970.073 | 4.9240.072 | 13.3440.051 | 10.5400.159 |
| +Non-st+SFL | MAE | 0.6710.006 | 4.8820.032 | 5.1760.054 | 5.0080.056 | 4.6530.036 | 11.9290.054 | 9.5390.054 |
| +Dish-TS | MAE | 0.7640.003 | 5.0240.080 | 5.3700.101 | 5.2150.017 | 4.8430.022 | 12.3630.314 | 9.8680.223 |
| +Dish-TS+SFL | MAE | 0.6760.008 | 4.9980.094 | 5.3170.079 | 5.0770.050 | 4.7770.053 | 12.2080.208 | 9.7200.133 |
| +ST-norm | MAE | 1.2880.038 | 5.2050.069 | 5.4690.072 | 9.7820.526 | 9.4550.500 | 13.9930.621 | 11.2520.529 |
| +T-norm | MAE | 1.7120.046 | 5.2910.245 | 5.5560.146 | 8.8760.935 | 8.7040.932 | 13.4690.292 | 10.7880.319 |
| +T-norm+SFL | MAE | 0.7470.165 | 4.9520.056 | 5.2390.031 | 8.4790.211 | 8.2040.240 | 12.1490.124 | 9.6540.055 |
V-D Ablation Studies
To evaluate the effectiveness of each module in our network, we conduct an ablation study by systematically removing specific components. Specifically, we remove the DSFL module, the gate mechanism in DSFL module, the SFL module, the entire distribution adaptive module, and the Stochastic Learner to observe the resulting changes in prediction accuracy. The ablation strategies are detailed as follows:
w/o Sto: We remove the Stochastic Learner and used only the deterministic features for prediction. In this case, only is input into the decoder.
w/o Sta: We remove all modules related to non-stationarity and distribution adaptation, including normalization and de-normalization operations and the SFL module, and replace the de-stationary attention with standard attention [64].
w/o SFL: We remove the SFL module, resulting in a temporal-only normalization similar to that in the non-stationary Transformer.
w/o DSFL: We remove the DSFL module and use spatial attention instead.
w/o Gate: We remove the gate mechanism and replace it with a Linear layer mapping concatenated feature to shape .
The results of the ablation study are depicted in Fig. 9. The final model, incorporating all modules, achieves the best performance. It is evident that the SFL module contributes most significantly to improving prediction accuracy across different tasks. Removing all distribution adaptation modules (w/o Sta) has less impact compared to removing the SFL module (w/o SFL), highlighting SFL as a crucial and indispensable component for spatio-temporal distribution adaptation. The effectiveness of the Stochastic Learner varies depending on the task, aligning with the fact that task uncertainties differ. Comparing the experimental results of "w/o Gate" and "w/o DSFL," the prediction error increases more significantly when the gate mechanism in the DSFL module is removed than when the entire DSFL module is removed. This suggests that fusing dynamic and static representations with a fixed ratio, without considering the time-varying relationships between them, hinders accurate prediction and results in an unsuitable feature combination. Moreover, our model outperforms the scenario where all normalization operations are removed, highlighting the necessity of normalization in spatio-temporal data processing.
VI Discussion and Conclusions
From the carried out experiments we can conclude the following:
The proposed method performs better in the spatio-temporal forecasting task compared with baseline methods at the cost of a moderate computation.
The proposed SFL module can be inserted with other temporal normalization methods and architectures, adapting distribution shifts in spatio-temporal context.
The proposed DSFL module is effective to both capture the dynamic and static spatial relations.
Each component (SFL, DSFL and Stochatic learner) in our DRAN model is provides value in prediction accuracy improvements. The adaptive fusion ratio derived from the gate mechanism is important for the integration of static and dynamic features.
Despite its strengths in accuracy performance, the methods suffers in terms of memory requirements and computational resources needs from scalability in large datasets or applications characterized by real-world systems. Additionally, the current design focuses on regular spatio-temporal patterns, making it less effective in scenarios characterized by the presence of abrupt changes or rare events.
To overcome above limitations, future work should focus on:
Scalability: Given the time costs associated with model training and inference, further research should investigate a more lightweight framework to deal with distribution shifts. This approach would improve scalability and enable the model to be effectively applied to large real-world datasets.
Adaptability and Transferability: While our framework currently emphasizes relation and distribution adaptation, it lacks mechanisms for dynamically adjusting network parameters based on learned knowledge and new inputs. Future work will focus on developing strategies to learn and update network parameters, drawing inspiration from techniques like EAST-Net [52], which generates sequence-specific, on-the-fly parameters. Enhancing adaptability will enable the model to better detect and respond to sudden changes and events.
VII Conclusion
To conclude, Spatio-temporal forecasting is essential for understanding the states of complex systems, yet accurate predictions are often hindered by the dynamic and intricate nature of these systems. This study addresses the challenge of adapting to dynamic changes in spatio-temporal systems using neural networks. We propose a DRAN to accommodate changes in distribution shifts, relations, and stochastic variations. Our approach includes a SFL to enable effective temporal normalization for spatio-temporal contexts. Additionally, we develop a DSFL to capture features from both dynamic and static relations. Furthermore, our framework enables to learn the deterministic and stochastic representations of features. Experimental results demonstrate the superiority of our method and the effectiveness of its components.

### v2

@@
 Abstract
-Accurate predictions of spatio-temporal systems’ states are crucial for tasks such as system management, control, and crisis prevention.
-However, the inherent time variance of spatio-temporal systems poses challenges to achieving accurate predictions whenever stationarity is not granted.
-To address non stationarity frameworks, we propose a Distribution and Relation Adaptive Network (DRAN) capable of dynamically adapting to relation and distribution changes over time.
+Accurate predictions of spatio-temporal systems are crucial for tasks such as system management, control, and crisis prevention.
+However, the inherent time variance of many spatio-temporal systems poses challenges to achieving accurate predictions whenever stationarity is not granted.
+In order to address non-stationarity, we propose a Distribution and Relation Adaptive Network (DRAN) capable of dynamically adapting to relation and distribution changes over time.
 While temporal normalization and de-normalization are frequently used techniques to adapt to distribution shifts, this operation is not suitable for the spatio-temporal context as temporal normalization scales the time series of nodes and possibly disrupts the spatial relations among nodes.
-In order to address this problem, we develop a Spatial Factor Learner (SFL) module that enables the normalization and de-normalization process in spatio-temporal systems.
+In order to address this problem, a Spatial Factor Learner (SFL) module is developed that enables the normalization and de-normalization process in spatio-temporal systems.
 To adapt to dynamic changes in spatial relationships among sensors, we propose a Dynamic-Static Fusion Learner (DSFL) module that effectively integrates features learned from both dynamic and static relations through an adaptive fusion ratio mechanism.
 Furthermore, we introduce a Stochastic Learner to capture the noisy components of spatio-temporal representations.
-Our approach outperforms state of the art methods in weather prediction and traffic flows forecasting tasks.
+Our approach outperforms state-of-the-art methods on weather prediction and traffic flow forecasting tasks.
 Experimental results show that our SFL efficiently preserves spatial relationships across various temporal normalization operations.
 Visualizations of the learned dynamic and static relations demonstrate that DSFL can capture both local and distant relationships between nodes.
-Moreover, ablation studies confirm the effectiveness of each component.
 Index Terms:
 Spatio-temporal forecasting, graph neural network, distribution adaptation, adaptive network.I Introduction
-Spatio-temporal systems, characterized by intricate spatial interactions among nodes and varying temporal dynamics, are prevalent in various fields such as physics [1], meteorology [2, 3, 4], power grids [5, 6] and transportation [7, 8, 9].
-Real world spatio-temporal systems often entail a large number of nodes, with interactions between nodes varying over time [10, 11].
-The complexity of these systems makes it challenging to make efficient decisions and manage future developments based on current conditions.
-Therefore, there is an urgent need for more accurate spatio-temporal prediction methods to support decision-making [12].
-In spatio-temporal forecasting, historical time series data associated with nodes are used to predict future observations of the same nodes [13, 14].
-Although numerous methods, particularly those based on deep learning have been developed [15, 16, 17, 18, 19], achieving accurate predictions remains challenging due to the time variance of the stochastic processes generating the data.
-This time variance can manifest in several ways, such as distribution shifts [20, 21], changes in spatial relationships among nodes [22], and changes in the nature of noise [23].
-Normalization and de-normalization mechanisms are commonly employed to address distribution shifts in time series forecasting [24, 25, 26, 27].
-Such techniques normalize input data to achieve consistent distributions and rebuild temporal distributions during de-normalization.
-Approaches such as Reversible Instance Normalization (RevIN) [24] and Dish-TS [27], which apply normalization and de-normalization on instance or temporal dimensions, have demonstrated their effectiveness in adapting to distribution shifts in time series forecasting.
-However, these methods may not be well-suited for spatio-temporal contexts.
-Methods that apply temporal normalization often scale the time series of nodes using different scaling factors.
-As a result, spatial message propagation occurs between nodes that have been scaled differently, which does not accurately reflect real-world message propagation processes.
+Spatio-temporal systems, characterized by intricate spatial interactions among senors (nodes) and rich temporal dynamics, are prevalent in various fields e.
+g., physics [1], meteorology [2, 3, 4], power grids [5, 6, 7] and transportation [8, 9].
+Such systems often entail a large number of nodes, with interactions among nodes varying over time [10, 11].
+The complexity of these systems and the time variance makes it challenging to efficiently manage future system developments and decision-making, hence requiring more accurate spatio-temporal prediction [12].
+In spatio-temporal forecasting, historical time series data associated with nodes are used to predict future observations [13, 14].
+Although numerous effective methods, particularly those based on deep learning, have been developed [15, 16, 17], predictions remains challenging due to the time variance of the environment-system interaction.
+This time variance can manifest in several ways, e.
+g., distribution shifts [18, 19], changes in spatial relationships among nodes [20], and changes in the nature of noise [21].
+Normalization and de-normalization mechanisms are commonly employed to address distribution shifts in time series forecasting, particularly in scenarios that show temporal variation without strong spatial interactions [22, 23, 24, 25].
+Normalization is applied to the input data to align its distribution with that of the training set, while de-normalization maps the processed results back to the original temporal distribution.
+However, above methods, which apply normalization and de-normalization on instance [22] or temporal dimensions [25], may not be well-suited for processing spatio-temporal data using graph neural networks that rely on relational inductive bias.
+In fact, temporal normalization scales the time series of nodes using different scaling factors as the temporal content evolves with time variance, hence causing erroneous scaled spatial messages propagating between nodes.
 Moreover, instance-level normalization uses the mean and standard deviation of all nodes over the input time span, addressing only coarse spatial and temporal distribution shifts, which is inadequate for adapting to the complex distributions in spatio-temporal contexts.
 Therefore, it is necessary to develop approaches that achieve temporal normalization while preserving spatial patterns in spatio-temporal forecasting tasks.
-To capture time-varying spatial relationships, some methods learn inherent static spatial relations from data [28, 29, 30], while others generate dynamic dependencies based on input windows [31, 32] .
-However, static relations remain fixed after training, limiting their adaptability, while dynamic relations derived from input data often exhibit instability in predictive performance.
-To address these limitations, some approaches jointly learn static and dynamic characteristics and fuse them with a weighted sum [33] or a fusion ratio constrained by loss functions [34], thus achieving more comprehensive spatial representation learning [35].
-However, these approaches fail to account for the dynamic interplay between static and dynamic components, relying on fixed fusion ratios that result in insufficient spatio-temporal representations.
-Therefore, there is a pressing need for a network capable of adapting to both static and dynamic relations while fusing their features using adaptive rations.
-Spatio-temporal systems inherently contain uncertainties due to data inaccuracies, noise, or sudden stochastic events within the systems.
+Various methods have been proposed to capture time-varying spatial relationships, including learning inherent static spatial structures from data [26, 27, 28], generating dynamic dependencies based on input windows [29, 30], and jointly modeling static and dynamic components through weighted fusion [31] or loss-constrained fusion ratios [32].
+However, these approaches often overlook the dynamic interplay between static and dynamic spatial components, relying on fixed fusion schemes.
+This can lead to suboptimal spatio-temporal representations, especially in scenarios where the relative importance of static and dynamic features varies over time.
+Therefore, there is a pressing need for methods capable of adapting to both static and dynamic relations.
+In addition, spatio-temporal systems are inherently affected by various sources of uncertainty, including data inaccuracies, sensor noise, and sudden stochastic events that might change with time.
 Given that these uncertainties impact prediction performance, various efforts have been made to quantify them.
@@
 Liang et al.
-[38] utilize latent variables to capture the uncertainties in air quality data, enhancing the accuracy of deterministic predictions.
+[35] utilize latent variables to capture the uncertainties in air quality data, enhancing the accuracy of predictions.
 In this paper, we propose a Distribution and Relation Adaptive Network (DRAN) that dynamically adapts to temporal variations and assesses uncertain components influenced by noise to achieve more comprehensive representation learning in spatio-temporal systems.
 More specifically, we develop a Spatial Factor Learner (SFL) that incorporates temporal normalization and de-normalization stages to mitigate the impact of distribution shifts while preserving spatial dependencies.
-To better accommodate spatio-temporal relations, we introduce a Dynamic-Static Fusion Learner (DSFL) module that fuses features learned from sample-specific dynamic and static relations using an adaptive fusion ratio mechanism.
-Additionally, we propose a Stochastic Learner based on a Variational Autoencoder (VAE) to estimate the noise.
-Experimental validations on weather and traffic systems to forecast temperature and traffic flow demonstrate that our model outperforms state-of-the-art methods.
-The SFL preserves the spatial distributions across various normalization methods.
-The learned dynamic and static relations capture non-overlapping local and distant node relations.
-Ablation studies further validate the effectiveness of each component.
-In summary, the novelty of our method is outlined as follows:
-We introduce a SFL module that enables a normalization and de-normalization mechanism for distribution adaptation in a spatio-temporal context.
+To better accommodate spatio-temporal relations, we introduce a Dynamic-Static Fusion Learner (DSFL) module that merges features learned from sample-specific dynamic and static relations using an adaptive fusion ratio mechanism.
+Additionally, we propose a Stochastic Learner based on a Variational Autoencoder (VAE) to estimate uncertainty.
+Experimental validations on weather and traffic systems demonstrate that our model outperforms state-of-the-art methods.
+The novelty of our method can be summarized as follows:
+We introduce an SFL module that enables a normalization and de-normalization mechanisms for distribution adaptation in a spatio-temporal context.
 The SFL module reverses temporal normalization while ensuring that spatial distributions closely match those of the forecasting horizons.
-Different from previous spatio-temporal relations adaptive methods which fuse dynamic and static features with a fixed ratio [35, 39, 33, 34], we propose a DSFL module to adaptively combine features from both dynamic and static relations with a gating mechanism, enabling accommodating the changing relations of dynamic and static perspectives.
-We propose a framework to learn the spatio-temporal representations from the deterministic and stochastic perspective.
-A VAE-based Stochastic Learner is introduced to learn the noisy components of representations.
-The remainder of this paper is organized as follows.
-Section II introduces the current work on learning spatial and temporal relations and adapting distributions.
+Different from previous spatio-temporal relations adaptive methods which fuse dynamic and static features with a fixed ratio [36, 37, 31, 32], we propose a DSFL module to adaptively combine features from both dynamic and static relations with a gating mechanism, accommodating the changing relations.
+We propose a framework to learn both deterministic and stochastic spatio-temporal representations, where a VAE-based Stochastic Learner is introduced to capture the noisy components of the representations.
+The paper is organized as follows.
+Section II introduces the state of the art.
 Section III details our network architecture and the overall workflow.
@@
 Section V presents the numerical results for traffic flow and weather predictions, accompanied by an ablation study and visualizations of module impacts.
-To have a general view of the experimental results and the effectiveness of modules, we make a discussion in Section VI.
-Finally, we conclude the paper and outline future research directions in Section VII.
+The discussion is provided in Section VI, and the conclusion in Section VII.
 II Related Works
-II-A Spatial and Temporal Relation Learning
-Spatio-temporal forecasting methods are designed to capture both spatial and temporal relationships from static and dynamic perspectives.
-Static relations are typically learned from training datasets using trainable node embeddings and adjacency matrices to depict inter-node relationships [40, 17, 41].
-Conversely, other studies [42, 43, 44] construct dynamic networks by generating sample-specific relations, which can be derived either directly from each input time series based on node similarity [45, 46] or through the use of meta-learners that generate dynamic parameters [47, 48].
+II-A Spatio-temporal forecasting
+Spatio-temporal forecasting is a longstanding and important research task.
+Early approaches primarily employed statistical models such as Historical Average (HA) and ARIMA [38, 39, 40], which offer interpretability but struggle to capture complex spatio-temporal dependencies.
+With the advent of deep learning, various modules are developed to model temporal and spatial dependencies more effectively.
+For temporal modeling, recurrent and convolutional neural networks [41, 42, 43] show effectiveness in short-term forecasting, while transformer-based architectures and state space models [44] demonstrate strong performance in both short- and long-term forecasting.
+Recurrent networks update historical information through hidden states, whereas transformer-based methods use attention mechanisms to capture dependencies from time-point [45], patch [46], and frequency-domain perspectives [47, 48].
+To model spatial dependencies, both explicit and implicit methods have been proposed.
+Explicit methods typically utilize graph-based structures to represent node-link relationships and facilitate message propagation [28].
+In contrast, implicit methods encode spatial dependencies within network parameters or attention mechanisms [49, 50, 51].
+Despite promising performance, many existing methods still struggle with capturing evolving spatial structures and often lack interpretability.
+To address these issues, decomposition-based approaches emerge as a promising direction.
+These methods decompose temporal signals into trend and seasonal components [52, 53], extract multi-frequency patterns [54], and learn multi-scale spatial relationships [55, 56].
+In this work, we propose a framework that decomposes spatial relations into dynamic and static components while learning both deterministic and stochastic features.
+II-B Static and dynamic Relation Learning
+Spatio-temporal forecasting methods are designed to capture both the static and dynamic relations existing among nodes.
+Static relations are typically learned from training datasets using trainable embeddings and adjacency matrices to depict inter-node relationships [57, 16, 58].
+Conversely, other studies [59, 60] construct dynamic networks by generating sample-specific relations, which can be derived either directly from each input time series based on node similarity [56, 61] or through the use of meta-learners that generate dynamic parameters [62, 63].
 Additionally, some methods refine the original relationships with input features [20, 64].
@@
 For instance, Fang et al.
-[34] decompose spatial relations into static and dynamic components, using a min-max learning paradigm with hyper-parameters to simultaneously learn both aspects.
+[32] decompose spatial relations into static and dynamic components, using a min-max learning paradigm with hyperparameters to simultaneously learn both aspects.
 Similarly, Li et al.
 [31] employ two parallel Graph Convolutional Networks (GCNs) to learn static and dynamic information, which is then fused using a weighted sum.
-However, these methods typically fuse the static and dynamic features at a fixed ratio.
-II-B Adaptation to Distribution Shifts
+These methods typically fuse the static and dynamic features at a fixed ratio.
+In this work, we focus on the adaptive fusion of dynamic and static spatial relations.
+II-C Adaptation to Distribution Shifts
 Normalization methods utilize statistic properties to normalize and de-normalize time series, adapting to distribution shifts in systems.
@@
 Each observation at time step is the result of a stochastic process, drawn from a conditioned distribution , where denotes the observations from nodes at time step in the spatio-temporal framework.
-The probability distribution varies over time and differs across nodes, indicating that distribution shifts occur when the observations across different time periods exhibit distinct distributions, i.e.,
+The probability distribution varies over time and differs across nodes, indicating that distribution shifts occur when observations across different time periods exhibit distinct distributions, i.e.,
 where is a small threshold, is a distance function that estimates the discrepancy between distributions, and and refer to two different time steps.
-In this work, we utilize Gaussian kernel density estimation [53] for distribution estimation and use KL divergence-as distance function.
+In this work, we utilize Gaussian kernel density estimation [68] for distribution estimation and use KL divergence as a distance function.
 The goal of the spatio-temporal forecasting task is to develop a model that uses the historical observations of length of nodes to predict the future -step observations of nodes , where refers to time series of observed variables of all the nodes ().
@@
 The distribution of historical time series often diverges from that of future time series due to time variance.
-Figure 2 (a) demonstrates the effectiveness of temporal normalization, which aligns the historical distribution more closely with the future distribution.
+Fig.
+2 (a) demonstrates the effectiveness of temporal normalization, which aligns the historical distribution more closely with the future distribution.
 This alignment facilitates easier learning of the projection from the input to the forecasting horizon.
@@
 Direct application of these stationary strategies to spatio-temporal forecasting results in degraded performance.
+As shown in Fig.
+2 (b), the direct application of normalization (colored in green) leads to spatial distributions deviating significantly from the spatial distribution of horizon windows.
 This is because they perform normalization within the time series of each node, leading to inconsistent scaling among nodes that are spatially interconnected.
 Therefore, we develop a framework that learns the temporal distribution shift on each node and preserves the spatial relations of nodes, thus maintaining the efficiency of spatial layers.
-To learn the deterministic part of the inputs, We firstly filter the time series by removing high-frequency noise.
+To learn the deterministic part of the inputs, the input time series is firstly filtered by removing high-frequency noise.
 We then normalize the input time series of each node using the temporal mean and standard deviation calculated for each input window.
@@
 III-C Dynamic-Static Fusion Learner
-Static neural networks with fixed parameters are inadequate for predicting dynamic systems.
-We consider that the relations of spatio-temporal systems consist of both static and dynamic components.
-Solely relying on input time series from different periods to establish spatio-temporal relations can lead to excessive fluctuations, thereby destabilizing prediction performance.
-Hence, it is essential to learn both dynamic and static information simultaneously and integrate them effectively.
+In this section, we propose the DSFL module, which simultaneously captures dynamic and static spatial information and integrates them through an adaptive fusion mechanism.
 To capture static information for a given task, we employ a trainable task-adaptive node embedding, akin to the approach used in the Adaptive Graph Convolutional Recurrent Network (AGCRN) [57].
@@
 where conducts concatenation operation on the last dimension of dynamic features and static features , generating the concatenated features .
-We then generate the gate signal by applying a Sigmoid activation function to ; FC denotes the fully connected layer.
+We generate the gate signal by applying a FC layer to map from to , followed by a Sigmoid activation function; here, FC denotes the fully connected layer.
 Finally, we employ Equation (12) to fuse the dynamic and static features, producing the final deterministic features , which represent the final deterministic features learned from the input time series.
@@
 The loss function consists of three components: prediction error , reconstruction error , and the distribution error computed using KL divergence .
-where and are hyper-parameters that balance the importance of various loss functions, and represents the time series generated by the neural network.
+where and are hyperparameters that balance the importance of various loss functions, and represents the time series generated by the neural network.
 Both the prediction error and the reconstruction error are computed using the Mean Absolute Error (MAE).
@@
 | Ours | ✓ | ✓ |
+| Weather | NYCBike1 | NYCBike2 | NYCTaxi | PeMS04 | PeMS08 | ||||||
+| MAE | MAE | MAE | MAE | MAE | MAE | ||||||
+| 0.30 | 0.765 | 1.50 | 5.115 | 1.00 | 4.903 | 1.50 | 10.961 | 0.50 | 19.283 | 0.50 | 13.884 |
+| 0.50 | 0.719 | 3.50 | 5.073 | 1.50 | 4.875 | 2.50 | 11.039 | 0.70 | 18.614 | 0.70 | 13.879 |
+| 0.80 | 0.705 | 5.50 | 5.158 | 2.50 | 4.914 | 3.50 | 10.864 | 0.90 | 18.585 | 0.90 | 13.933 |
+| 1.00 | 0.660 | 7.50 | 5.057 | 3.50 | 4.827 | 4.50 | 11.074 | 1.00 | 18.561 | 1.00 | 13.726 |
+| 3.00 | 0.766 | 9.50 | 5.159 | 4.50 | 4.847 | 5.50 | 11.030 | 2.00 | 18.615 | 1.20 | 13.696 |
+| MAE | MAE | MAE | MAE | MAE | MAE | ||||||
+| 0.10 | 0.670 | 0.10 | 5.390 | 0.10 | 5.130 | 0.10 | 11.264 | 0.10 | 32.290 | 0.10 | 20.790 |
+| 0.50 | 0.665 | 0.50 | 5.269 | 0.50 | 4.910 | 0.50 | 10.758 | 0.50 | 25.774 | 0.50 | 13.818 |
+| 1.00 | 0.664 | 1.00 | 5.173 | 1.00 | 4.866 | 1.00 | 11.136 | 1.00 | 18.585 | 1.00 | 13.717 |
+| 5.00 | 0.658 | 5.00 | 4.982 | 5.00 | 5.025 | 5.00 | 11.114 | 5.00 | 18.545 | 5.00 | 13.822 |
+| 10.00 | 0.672 | 10.00 | 5.303 | 10.00 | 4.938 | 10.00 | 11.161 | 10.00 | 18.387 | 10.00 | 13.696 |
+| Selection () | (1.00, 5.00) | () | (7.50, 5.00) | () | (3.50, 1.00) | () | (3.50, 0.50) | () | (1.00,10.00) | () | (1.00, 1.00) |
 After obtaining the deterministic component and the stochastic component of the time series, we input them into the decoder to map the features to the forecasting results.
@@
 IV Experimental Setups
-In this section, we detail our experimental setup, including the datasets used, the hyper-parameter settings for our networks, the baseline comparisons, and the training process.
+In this section, we detail our experimental setup, including the datasets used, the hyperparameter settings for our networks, the baseline comparisons, and the training process.
 IV-A Datasets
@@
 Application features of these datasets are shown in Table I.
-| Weather | NYCBike1 | NYCBike2 | NYCTaxi | PeMS04 | PeMS08 | ||||||
-| MAE | MAE | MAE | MAE | MAE | MAE | ||||||
-| 0.30 | 0.765 | 1.50 | 5.115 | 1.00 | 4.903 | 1.50 | 10.961 | 0.50 | 19.283 | 0.50 | 13.884 |
-| 0.50 | 0.719 | 3.50 | 5.073 | 1.50 | 4.875 | 2.50 | 11.039 | 0.70 | 18.614 | 0.70 | 13.879 |
-| 0.80 | 0.705 | 5.50 | 5.158 | 2.50 | 4.914 | 3.50 | 10.864 | 0.90 | 18.585 | 0.90 | 13.933 |
-| 1.00 | 0.660 | 7.50 | 5.057 | 3.50 | 4.827 | 4.50 | 11.074 | 1.00 | 18.561 | 1.00 | 13.726 |
-| 3.00 | 0.766 | 9.50 | 5.159 | 4.50 | 4.847 | 5.50 | 11.030 | 2.00 | 18.615 | 1.20 | 13.696 |
-| MAE | MAE | MAE | MAE | MAE | MAE | ||||||
-| 0.10 | 0.670 | 0.10 | 5.390 | 0.10 | 5.130 | 0.10 | 11.264 | 0.10 | 32.290 | 0.10 | 20.790 |
-| 0.50 | 0.665 | 0.50 | 5.269 | 0.50 | 4.910 | 0.50 | 10.758 | 0.50 | 25.774 | 0.50 | 13.818 |
-| 1.00 | 0.664 | 1.00 | 5.173 | 1.00 | 4.866 | 1.00 | 11.136 | 1.00 | 18.585 | 1.00 | 13.717 |
-| 5.00 | 0.658 | 5.00 | 4.982 | 5.00 | 5.025 | 5.00 | 11.114 | 5.00 | 18.545 | 5.00 | 13.822 |
-| 10.00 | 0.672 | 10.00 | 5.303 | 10.00 | 4.938 | 10.00 | 11.161 | 10.00 | 18.387 | 10.00 | 13.696 |
-| Selection () | (1.00, 5.00) | () | (7.50, 5.00) | () | (3.50, 1.00) | () | (3.50, 0.50) | () | (1.00,10.00) | () | (1.00, 1.00) |
 IV-B Training Details
@@
 The decoder comprises 2 Linear layers that fuse the deterministic and stochastic representations to produce the target feature outputs.
-The balance hyper-parameters and are fine-tuned experimentally to account for the stochastic nature and uncertainties of the datasets.
+The balance hyperparameters and are fine-tuned experimentally to account for the stochastic nature and uncertainties of the datasets.
 Table III presents the variation in prediction error corresponding to different values of , with the value minimizing the error selected as optimal.
@@
 ASTGCN and STAEformer employ attention mechanisms to capture dynamic changes in input features.
-RGSL learns spatio-temporal dependencies from a predefined graph and learnable node embeddings.
-It dynamically fuses features from two graphs using an attention mechanism.
+RGSL learns spatio-temporal dependencies from a predefined graph and learnable node embeddings, dynamically fusing features from two graphs using an attention mechanism.
 Additionally, STAEformer employs learnable node embeddings and concatenates it with input time series to capture static features.
@@
 TESTAM employs a mixture-of-experts model with three experts: one for temporal modeling, one for spatio-temporal modeling with a static graph, and one for spatio-temporal dependency modeling with a dynamic graph.
+We also fine-tune the hyperparameters of baseline methods to ensure fairness, especially for datasets not included in their official implementations.
 We evaluate the performance of these models using two metrics: MAE and Mean Absolute Percentage Error (MAPE).
@@
 V-A Comparison Results
-Numerical experiments, as shown in Tables IV and V, present the average performance and standard deviation of both the baseline methods and our DRAN model.
+As shown in Tables IV and V, numerical experiments, present the average performance and standard deviation of both the baseline methods and our DRAN model.
 Our DRAN model outperforms the baseline methods, demonstrating the best overall performance.
-Specifically, we observe a reduction in MAE of up to 7.8% in weather forecasting and 1.7% in traffic flow prediction compared to the baselines.
-DRAN achieves the best average performance on the Weather and NYC datasets.
+DRAN achieves the best average performance on the Weather and NYC datasets, and ranks second on the PeMS04 and PeMS08 datasets.
+Our method performs less effectively on the MAPE metric but excels on the MAE metric.
+This suggests that the model prioritizes optimizing the overall error rather than minimizing relative error in specific scenarios, such as scenarios with small target values.
+Consequently, while DRAN’s performance on the MAPE metric is slightly lower than that of other methods, its superior MAE results demonstrate effective overall error control.
 In terms of prediction stability, DRAN exhibits slightly larger standard deviations compared to other methods across experiments with different random seeds.
 This is likely due to the model’s attempt to learn noise components in the representation, which are randomly sampled from distributions.
-Our method performs less effectively on the MAPE metric but excels on the MAE metric.
-This suggests that the model prioritizes optimizing overall error rather than minimizing relative error in specific scenarios, such as scenarios with small target values.
-Consequently, while DRAN’s performance on the MAPE metric is slightly lower than that of other methods, its superior MAE results demonstrate effective overall error control.
 ST-SSL exhibits strong performance in one-step traffic flow prediction but is less effective on other datasets.
@@
 | Temperature | Inflow | Outflow | Inflow | Outflow | Inflow | Outflow | ||
-| DA-RNN [54] | MAE | 7.9511.036 | 14.1583.523 | 14.5653.383 | 13.4461.066 | 13.5131.382 | 72.88012.173 | 56.05610.753 |
-| MAPE (%) | 2.7090.355 | 60.35912.515 | 62.95512.885 | 58.67813.627 | 58.80714.098 | 107.44121.603 | 98.41417.457 | |
+| DA-RNN [69] | MAE | 7.9511.036 | 14.1583.523 | 14.5653.383 | 13.1921.014 | 13.1151.178 | 24.6427.398 | 28.7218.953 |
+| MAPE (%) | 2.7090.355 | 60.35912.515 | 62.95512.885 | 53.8944.686 | 60.25313.219 | 67.3199.888 | 70.73714.281 | |
 | InfoTS [70] | MAE | 1.3430.031 | 6.4360.133 | 6.7970.138 | 6.6410.091 | 6.0870.096 | 14.9120.239 | 11.5450.225 |
@@
 | MAPE (%) | 0.7780.377 | 58.2081.926 | 58.7891.383 | 63.40219.134 | 62.23713.177 | 84.12430.571 | 75.09424.967 | |
-| TGCN [28] | MAE | 1.7360.847 | 7.4120.052 | 7.6770.441 | 12.18510.559 | 10.7928.433 | 25.30711.321 | 21.7159.109 |
-| MAPE (%) | 0.5920.287 | 34.6110.827 | 35.0851.704 | 37.5328.401 | 36.0457.883 | 44.53610.924 | 45.18110.584 | |
-| MemDA [61] | MAE | 1.7860.071 | 6.7770.154 | 7.2460.100 | 6.7800.257 | 6.3440.286 | 20.9391.515 | 20.6311.268 |
-| MAPE (%) | 0.6110.026 | 29.3440.543 | 30.3510.126 | 30.0990.342 | 28.5391.064 | 29.6513.243 | 34.9792.950 | |
-| ASTGCN [57] | MAE | 1.5210.155 | 6.4810.352 | 6.6980.541 | 8.7235.180 | 7.5483.533 | 15.3895.711 | 12.5003.923 |
-| MAPE (%) | 0.5200.054 | 30.7801.221 | 31.2641.976 | 28.6771.634 | 27.1992.117 | 24.7310.916 | 24.4431.578 | |
-| TESTAM [60] | MAE | 1.4811.203 | 6.3310.736 | 6.8260.891 | 6.5580.941 | 6.3081.003 | 22.6863.816 | 21.4504.195 |
-| MAPE (%) | 0.5070.415 | 30.4182.571 | 31.6282.949 | 30.1603.249 | 29.7054.047 | 32.1435.566 | 37.2377.208 | |
-| AGCRN [40] | MAE | 4.8522.311 | 6.3220.847 | 6.5250.510 | 10.8645.159 | 9.8893.836 | 19.7868.580 | 16.4326.402 |
-| MAPE (%) | 1.6690.798 | 30.0491.915 | 30.5751.184 | 34.9993.415 | 34.0323.307 | 33.3076.145 | 32.1815.601 | |
-| GCGRU [30] | MAE | 1.0120.041 | 5.4570.063 | 5.6310.290 | 7.0363.489 | 6.2502.526 | 12.1532.649 | 10.2771.340 |
-| MAPE (%) | 0.3460.015 | 26.8670.700 | 27.2921.476 | 24.3372.990 | 23.7112.231 | 22.3867.043 | 22.8007.607 | |
-| DCRNN [58] | MAE | 0.9840.028 | 5.3740.039 | 5.5570.305 | 6.9403.500 | 6.1532.513 | 12.0973.639 | 9.8332.314 |
-| MAPE (%) | 0.3360.010 | 26.7720.850 | 27.1651.507 | 24.0703.034 | 23.5652.251 | 21.1823.214 | 21.5983.360 | |
+| TGCN [26] | MAE | 1.7360.847 | 7.4120.052 | 7.6770.441 | 6.1030.106 | 5.7120.072 | 25.30711.321 | 21.7159.109 |
+| MAPE (%) | 0.5920.287 | 34.6110.827 | 35.0851.704 | 29.7170.592 | 28.2860.632 | 44.53610.924 | 45.18110.584 | |
+| MemDA [76] | MAE | 1.7860.071 | 6.7770.154 | 7.2460.100 | 6.4520.127 | 6.2840.252 | 18.5050.661 | 17.1990.542 |
+| MAPE (%) | 0.6110.026 | 29.3440.543 | 30.3510.126 | 29.3120.571 | 27.9950.802 | 28.7191.861 | 32.4631.726 | |
+| ASTGCN [72] | MAE | 1.5210.155 | 6.4810.352 | 6.6980.541 | 6.6580.326 | 7.5483.533 | 15.3895.711 | 12.5003.923 |
+| MAPE (%) | 0.5200.054 | 30.7801.221 | 31.2641.976 | 28.6741.021 | 27.0281.123 | 24.7310.916 | 24.4431.578 | |
+| TESTAM [75] | MAE | 1.4811.203 | 6.3310.736 | 6.8260.891 | 6.4370.852 | 6.2660.901 | 17.6841.814 | 15.9672.758 |
+| MAPE (%) | 0.5070.415 | 30.4182.571 | 31.6282.949 | 28.5653.457 | 27.7622.620 | 25.7033.433 | 25.9884.105 | |
+| AGCRN [57] | MAE | 1.0940.039 | 5.0660.020 | 5.3660.023 | 5.2790.058 | 4.9410.035 | 12.2630.132 | 9.7200.098 |
+| MAPE (%) | 0.3730.013 | 25.5510.200 | 26.5640.209 | 25.4980.268 | 24.6330.286 | 18.2340.324 | 18.2390.278 | |
+| GCGRU [28] | MAE | 1.0120.041 | 5.4570.063 | 5.6310.290 | 5.3130.064 | 4.9870.056 | 12.1532.649 | 10.2771.340 |
+| MAPE (%) | 0.3460.015 | 26.8670.700 | 27.2921.476 | 25.4280.423 | 24.4900.185 | 22.3867.043 | 22.8007.607 | |
+| DCRNN [73] | MAE | 0.9840.028 | 5.3740.039 | 5.5570.305 | 5.3860.036 | 5.0470.031 | 12.0973.639 | 9.8332.314 |
+| MAPE (%) | 0.3360.010 | 26.7720.850 | 27.1651.507 | 25.2790.273 | 24.3320.164 | 21.1823.214 | 21.5983.360 | |
 | STAEformer [30] | MAE | 3.7282.367 | 5.1680.029 | 5.4750.028 | 5.4530.100 | 5.1120.147 | 12.2620.237 | 9.8240.109 |
 | MAPE (%) | 1.2840.819 | 25.8290.331 | 26.8890.242 | 25.7740.429 | 24.8400.447 | 17.8890.777 | 18.2030.684 | |
-| MegaCRN [59] | MAE | 0.9520.027 | 5.0420.016 | 5.3570.043 | 6.6023.125 | 5.8782.164 | 12.2060.083 | 9.7400.074 |
-| MAPE (%) | 0.3250.010 | 25.3290.167 | 26.2750.213 | 23.4233.445 | 22.7102.775 | 18.0310.582 | 17.9720.294 | |
-| RGSL [17] | MAE | 0.7270.003 | 5.1490.140 | 5.3350.200 | 6.9213.513 | 6.0822.518 | 13.9451.775 | 11.8592.918 |
-| MAPE (%) | 0.2480.001 | 25.6120.183 | 26.1100.924 | 24.1862.957 | 23.2882.249 | 27.14817.848 | 27.37117.905 | |
-| ST-SSL [22] | MAE | 1.3940.040 | 5.1350.024 | 5.2650.023 | 5.0420.029 | 4.7140.023 | 12.0100.481 | 9.7900.101 |
-| MAPE (%) | 0.4750.013 | 25.4300.295 | 24.6050.265 | 22.6330.112 | 21.8130.808 | 16.3830.100 | 16.8550.228 | |
+| MegaCRN [74] | MAE | 0.9520.027 | 5.0420.016 | 5.3570.043 | 5.1520.049 | 4.8260.021 | 12.0840.158 | 9.5690.175 |
+| MAPE (%) | 0.3250.010 | 25.3290.167 | 26.2750.213 | 24.8780.409 | 23.8700.293 | 17.8830.204 | 17.6640.338 | |
+| RGSL [16] | MAE | 0.7270.003 | 5.1490.140 | 5.3350.200 | 5.2510.039 | 4.9150.060 | 12.0830.125 | 9.5820.072 |
+| MAPE (%) | 0.2480.001 | 25.6120.183 | 26.1100.924 | 24.1862.957 | 23.2882.249 | 18.4300.531 | 18.2050.319 | |
+| ST-SSL [20] | MAE | 1.3520.010 | 5.1350.024 | 5.2650.023 | 5.0420.029 | 4.7140.023 | 12.0100.481 | 9.7900.101 |
+| MAPE (%) | 0.4610.004 | 25.4300.295 | 24.6050.265 | 22.6330.112 | 21.8130.808 | 16.3830.100 | 16.8550.228 | |
 | Ours | MAE | 0.6720.007 | 4.8820.032 | 5.1760.054 | 5.0080.056 | 4.6530.036 | 11.9290.054 | 9.5390.054 |
@@
 6, significant discrepancies between the spatial distribution of representations before SFL and the final horizon are observed.
-The SFL module effectively reduces these variations in the spatial distribution of learned representations, thereby facilitating the learning process to map spatially preserved representations to predictions.
+The SFL module effectively reduces these variations in the spatial distributions of learned representations, thereby facilitating the learning process to map spatially preserved representations to predictions.
+| PeMS04 | PeMS08 | ||
+| Flow | Flow | ||
+| DA-RNN [69] | MAE | 104.74112.184 | 92.2769.643 |
+| MAPE (%) | 137.64018.494 | 66.77413.187 | |
+| InfoTS [70] | MAE | 25.8510.510 | 24.0061.225 |
+| MAPE (%) | 19.5560.464 | 14.6820.666 | |
+| AutoTCL [71] | MAE | 23.8140.048 | 17.3550.200 |
+| MAPE (%) | 20.8790.098 | 13.0060.121 | |
+| TGCN [26] | MAE | 34.7850.204 | 24.4480.129 |
+| MAPE (%) | 27.9720.754 | 14.5890.219 | |
+| GCGRU [28] | MAE | 24.3980.043 | 17.7783.356 |
+| MAPE (%) | 15.1050.464 | 10.6222.838 | |
+| STGCN [27] | MAE | 25.3022.316 | 25.33810.825 |
+| MAPE (%) | 24.3218.603 | 15.4047.083 | |
+| DCRNN [73] | MAE | 24.9122.064 | 20.8530.062 |
+| MAPE (%) | 17.7200.973 | 11.8640.094 | |
+| ASTGCN [72] | MAE | 23.5650.942 | 20.1950.398 |
+| MAPE (%) | 16.8131.004 | 11.3030.552 | |
+| ST-SSL [20] | MAE | 22.8360.705 | 18.5590.097 |
+| MAPE (%) | 14.3180.435 | 10.6150.122 | |
+| MemDA [76] | MAE | 20.0370.150 | 16.3700.207 |
+| MAPE (%) | 11.9690.100 | 9.3420.243 | |
+| RGSL [16] | MAE | 19.5440.2571 | 17.4523.102 |
+| MAPE (%) | 13.8620.242 | 9.1860.182 | |
+| TESTAM [75] | MAE | 19.3310.481 | 15.7570.360 |
+| MAPE (%) | 12.0980.419 | 9.0830.215 | |
+| AGCRN [57] | MAE | 19.32910.3053 | 16.1010.085 |
+| MAPE (%) | 12.9370.035 | 9.2190.065 | |
+| MegaCRN [74] | MAE | 18.8580.0413 | 15.5970.245 |
+| MAPE (%) | 12.8080.098 | 8.8340.096 | |
+| STAEformer [30] | MAE | 18.2410.082 | 13.5380.039 |
+| MAPE (%) | 12.0640.071 | 8.8580.017 | |
+| Ours | MAE | 18.3750.087 | 13.6900.085 |
+| MAPE (%) | 12.0260.419 | 8.9870.057 | |
+Results with bold are the overall best performance, and shading results have the suboptimal performance.
+| Methods | Time (s) | MAE |
+| GConvGRU [28] | 1.383 | 12.022 |
+| DCRNN [73] | 0.967 | 11.924 |
+| TGCN [26] | 0.469 | 28.475 |
+| STGCN [27] | 0.343 | 12.286 |
+| RGSL [16] | 0.246 | 10.833 |
+| InfoTS [70] | 0.179 | 13.228 |
+| MegaRCN [74] | 0.178 | 10.827 |
+| AGCRN [57] | 0.126 | 10.992 |
+| ASTGCN [72] | 0.114 | 14.422 |
+| TESTAM [75] | 0.075 | 16.826 |
+| DRAN (Ours) | 0.075 | 10.734 |
+| ST-SSL [20] | 0.073 | 10.996 |
+| STAEformer [30] | 0.065 | 10.810 |
+| MemDA [76] | 0.056 | 17.852 |
+| AutoTCL [71] | 0.052 | 13.119 |
+| DA-RNN [69] | 0.048 | 26.596 |
+Experiments are conducted on a batch of NYCTaxi date which contains 8 samples. Results with bold are the fastest method and the method with best prediction performance.
+| Weather | NYCBike1 | NYCBike2 | NYCTaxi | |||||
+| Temperature | Inflow | Outflow | Inflow | Outflow | Inflow | Outflow | ||
+| +RevIN | MAE | 0.7320.004 | 5.0310.064 | 5.3410.094 | 5.1920.068 | 4.8310.060 | 12.3010.217 | 9.7140.168 |
+| +DAIN | MAE | 1.0350.006 | 5.2120.088 | 5.5080.084 | 5.6700.420 | 5.2580.290 | 13.7470.174 | 10.8480.088 |
+| +DAIN+SFL | MAE | 0.6630.003 | 4.9420.050 | 5.2290.028 | 5.2680.135 | 4.8990.159 | 12.3030.147 | 9.8280.171 |
+| +Non-st | MAE | 0.7380.006 | 5.0960.109 | 5.3680.089 | 5.2970.073 | 4.9240.072 | 13.3440.051 | 10.5400.159 |
+| +Non-st+SFL | MAE | 0.6710.006 | 4.8820.032 | 5.1760.054 | 5.0080.056 | 4.6530.036 | 11.9290.054 | 9.5390.054 |
+| +Dish-TS | MAE | 0.7640.003 | 5.0240.080 | 5.3700.101 | 5.2150.017 | 4.8430.022 | 12.3630.314 | 9.8680.223 |
+| +Dish-TS+SFL | MAE | 0.6760.008 | 4.9980.094 | 5.3170.079 | 5.0770.050 | 4.7770.053 | 12.2080.208 | 9.7200.133 |
+| +ST-norm | MAE | 1.2880.038 | 5.2050.069 | 5.4690.072 | 9.7820.526 | 9.4550.500 | 13.9930.621 | 11.2520.529 |
+| +T-norm | MAE | 1.7120.046 | 5.2910.245 | 5.5560.146 | 8.8760.935 | 8.7040.932 | 13.4690.292 | 10.7880.319 |
+| +T-norm+SFL | MAE | 0.7470.165 | 4.9520.056 | 5.2390.031 | 8.4790.211 | 8.2040.240 | 12.1490.124 | 9.6540.055 |
 Furthermore, we replace the temporal normalization modules to verify the effectiveness of SFL across various temporal normalization operations.
@@
 The dynamic adjacency matrix highlights the strength of relationships between nodes with similar signal patterns, whereas the static adjacency matrix focuses on the signals of individual nodes and some distributed nodes within the network.
-The differing concentrations of dynamic and static perspectives allow the model to learn features from various aspects.
+The contrast in focus enables the model to learn complementary features from both dynamic and static perspectives.
 To clarify the differences between the learned dynamic and static relations for specific nodes, we select node from various locations and visualize the relationships between the selected node and other nodes, represented as and .
@@
 This demonstrates that DSFL effectively learns comprehensive and non-overlapping spatial relations.
-| PeMS04 | PeMS08 | ||
-| Flow | Flow | ||
-| DA-RNN [54] | MAE | 130.38427.184 | 110.05619.332 |
-| MAPE (%) | 178.98325.358 | 100.13928.297 | |
-| InfoTS [55] | MAE | 25.8510.510 | 24.0061.225 |
-| MAPE (%) | 19.5560.464 | 14.6820.666 | |
-| AutoTCL [56] | MAE | 23.8140.048 | 17.3550.200 |
-| MAPE (%) | 20.8790.098 | 13.0060.121 | |
-| TGCN [28] | MAE | 34.78590.2043 | 33.6049.765 |
-| MAPE (%) | 27.9720.754 | 26.1729.541 | |
-| GCGRU [30] | MAE | 25.83360.0399 | 22.2658.494 |
-| MAPE (%) | 17.7880.442 | 14.3018.571 | |
-| STGCN [29] | MAE | 25.30172.3161 | 25.33810.825 |
-| MAPE (%) | 24.3218.603 | 15.4047.083 | |
-| DCRNN [58] | MAE | 24.91172.0638 | 20.8530.062 |
-| MAPE (%) | 17.7200.973 | 11.8640.094 | |
-| ASTGCN [57] | MAE | 23.56480.9421 | 20.3080.967 |
-| MAPE (%) | 16.8131.004 | 11.3030.552 | |
-| ST-SSL [22] | MAE | 23.1461.074 | 18.9890.713 |
-| MAPE (%) | 14.4130.697 | 10.7980.332 | |
-| MemDA [61] | MAE | 20.0370.150 | 16.3700.207 |
-| MAPE (%) | 11.9690.100 | 9.3420.243 | |
-| RGSL [17] | MAE | 19.544+-0.2571 | 17.4523.102 |
-| MAPE (%) | 13.8620.242 | 9.1860.182 | |
-| TESTAM [60] | MAE | 19.3310.481 | 15.7570.360 |
-| MAPE (%) | 12.0980.419 | 9.0830.215 | |
-| AGCRN [40] | MAE | 19.32910.3053 | 17.7902.210 |
-| MAPE (%) | 12.9370.035 | 10.3241.078 | |
-| MegaCRN [59] | MAE | 18.8580.0413 | 15.5970.245 |
-| MAPE (%) | 12.8080.098 | 8.8340.096 | |
-| STAEformer [32] | MAE | 18.2410.082 | 13.5380.039 |
-| MAPE (%) | 12.0640.071 | 8.8580.017 | |
-| Ours | MAE | 18.3750.087 | 13.6900.085 |
-| MAPE (%) | 12.0260.419 | 8.9870.057 | |
-Results with bold are the overall best performance, and shading results have the suboptimal performance.
-| Methods | Time (s) | MAE |
-| GConvGRU [30] | 1.383 | 12.022 |
-| DCRNN [58] | 0.967 | 11.924 |
-| TGCN [28] | 0.469 | 28.475 |
-| STGCN [29] | 0.343 | 12.286 |
-| RGSL [17] | 0.227 | 11.916 |
-| InfoTS [55] | 0.179 | 13.228 |
-| MegaRCN [59] | 0.166 | 10.970 |
-| AGCRN [40] | 0.134 | 18.338 |
-| ASTGCN [57] | 0.114 | 14.422 |
-| TESTAM [60] | 0.075 | 24.303 |
-| DRAN (Ours) | 0.075 | 10.737 |
-| ST-SSL [22] | 0.073 | 10.996 |
-| STAEformer [32] | 0.065 | 10.810 |
-| AutoTCL [56] | 0.052 | 13.119 |
-| MemDA [61] | 0.047 | 20.521 |
-| DA-RNN [54] | 0.038 | 64.468 |
-Experiments are conducted on a batch of NYCTaxi date which contains 8 samples. Results with bold are the fastest method and the method with best prediction performance.
-| Weather | NYCBike1 | NYCBike2 | NYCTaxi | |||||
-| Temperature | Inflow | Outflow | Inflow | Outflow | Inflow | Outflow | ||
-| +RevIN | MAE | 0.7320.004 | 5.0310.064 | 5.3410.094 | 5.1920.068 | 4.8310.060 | 12.3010.217 | 9.7140.168 |
-| +DAIN | MAE | 1.0350.006 | 5.2120.088 | 5.5080.084 | 5.6700.420 | 5.2580.290 | 13.7470.174 | 10.8480.088 |
-| +DAIN+SFL | MAE | 0.6630.003 | 4.9420.050 | 5.2290.028 | 5.2680.135 | 4.8990.159 | 12.3030.147 | 9.8280.171 |
-| +Non-st | MAE | 0.7380.006 | 5.0960.109 | 5.3680.089 | 5.2970.073 | 4.9240.072 | 13.3440.051 | 10.5400.159 |
-| +Non-st+SFL | MAE | 0.6710.006 | 4.8820.032 | 5.1760.054 | 5.0080.056 | 4.6530.036 | 11.9290.054 | 9.5390.054 |
-| +Dish-TS | MAE | 0.7640.003 | 5.0240.080 | 5.3700.101 | 5.2150.017 | 4.8430.022 | 12.3630.314 | 9.8680.223 |
-| +Dish-TS+SFL | MAE | 0.6760.008 | 4.9980.094 | 5.3170.079 | 5.0770.050 | 4.7770.053 | 12.2080.208 | 9.7200.133 |
-| +ST-norm | MAE | 1.2880.038 | 5.2050.069 | 5.4690.072 | 9.7820.526 | 9.4550.500 | 13.9930.621 | 11.2520.529 |
-| +T-norm | MAE | 1.7120.046 | 5.2910.245 | 5.5560.146 | 8.8760.935 | 8.7040.932 | 13.4690.292 | 10.7880.319 |
-| +T-norm+SFL | MAE | 0.7470.165 | 4.9520.056 | 5.2390.031 | 8.4790.211 | 8.2040.240 | 12.1490.124 | 9.6540.055 |
+Moreover, the gate mechanism enables adaptive integration of dynamic and static features.
+We visualize the first channel of the gate signal from three layers of DSFL modules in Fig.
+9, illustrating fusion ratio variations on the NYCBike2 test set.
+Each gate mechanism performs hierarchical fusion, where the blue-filled area represents the contribution of static features to the final deterministic features , and the orange area represents the dynamic features.
+The fusion ratios vary in sync with the fluctuations of the input inflow and outflow series, demonstrating the gate mechanism’s ability to adaptively balance dynamic and static information.
 V-D Ablation Studies
@@
 The results of the ablation study are depicted in Fig.
-9.
+10.
 The final model, incorporating all modules, achieves the best performance.
@@
 This study addresses the challenge of adapting to dynamic changes in spatio-temporal systems using neural networks.
-We propose a DRAN to accommodate changes in distribution shifts, relations, and stochastic variations.
+A DRAN framework is proposed to accommodate changes in distribution shifts, relations, and stochastic variations.
 Our approach includes a SFL to enable effective temporal normalization for spatio-temporal contexts.